HIGH-RESOLUTION NEAR-INFRARED SPECTROSCOPY OF H_{3}^{+} ABOVE THE BARRIER TO LINEAR-ITY

\underline{J. GOTTFRIED}, Department of Chemistry and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637; B. J. McCALL, Departments of Chemistry and Astronomy, University of California, Berkeley, CA 94720; T. OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, University of Chicago, Chicago, IL 60637.

Twenty transitions from the fourth overtone and combination bands of H_{3}^{+} (5ν_{2}^{1}, 5ν_{2}^{3}, 5ν_{2}^{5}, 2ν_{1} + 2ν_{2}^{2},…) have been observed. These transitions, which are more than 4600 times weaker than the fundamental band, occur in the near-infrared region and probe energy levels above the barrier to linearity (>10,000 cm^{−1}), the regime in which H_{3}^{+} has enough energy to sample linear configurations. The detection of these transitions required the development of a high-resolution, highsensitivity (ΔI/I∼1×10^{−7}) spectrometer based on a Ti:Sapphire laser and incorporating velocity modulation, heterodyne detection, noise subtraction, and multi-passing. Both pure hydrogen and He/H_{2} plasmas were used to discriminate between H_{3}^{+} and Rydberg transitions of H_{2}. The primary motivation for continuing the study of vibrational states beyond those spectroscopically probed to date^{a} is to assist in the development of theoretical calculations of H_{3}^{+}. The measured rovibrational energy levels provide an experimental check of ab initio calculations in this region, which present a unique challenge to theorists^{b}^{c}^{d}.

aC. M. Lindsay and B. J. McCall, J. Mol. Spectrosc. \underline{210}, 60 (2001).

bA. Alijah, P. Schiffels, and J. Hinze, manuscript in preparation, (2001).

cL. Neale, S. Miller, J. Tennyson, Astrophys. J. \underline{464}, 516, (1996).

dJ.K.G. Watson, personal communication, 1996.